### v1

@@ V-D2 Optimal Dimension for Short Queries: Headline Embeddings
The results, presented in Table X, reveal a distinct optimum for the headline embedding space. The 64-dimensional PCA-reduced embeddings achieved the highest overlap score of 63.1%, surpassing both 128D (62.4%) and 256D (61.4%). Figure  35 shows bar graph and heatmap analysis of statistics in table Table X.
TABLE X: Performance of PCA-reduced headline embeddings across dimensions at 100-character query length
Query Length Average Overlap with 768D (%) Optimal Dimension
(Characters) 64D 128D 256D
100 63.1 62.4 61.4
Mean 63.1 62.4 61.4
Figure 35: Vertical bar chart showing average overlap percentages for different reduced dimensions (64D, 128D, 256D) for headline embeddings
This outcome indicates that for the semantically dense and focused space of headlines, a lower-dimensional representation (64D) is not only sufficient but optimal. The concise nature of headlines, where core meaning is concentrated in fewer tokens, allows for effective semantic capture in a more compact vector space. Increasing the dimension to 128D or 256D provides no improvement and, in fact, leads to marginally decreased overlap, likely due to the introduction of redundant or noisy dimensions that do not contribute meaningful variance for this specific data distribution.
Consequently, 64D was selected as the optimal reduced dimension for headline-based retrieval. This choice aligns with the principle of matching representation complexity to data complexity: the simpler, more focused semantics of headlines are best preserved in a lower-dimensional space, ensuring efficient storage and retrieval without compromising the fidelity required for accurate short-query recommendations.
@@ V-E Final System Configuration and Performance Evaluation
Based on the comprehensive experimental analysis presented in above sections, the final retrieval pipeline of the ULTRA framework is configured as follows:
  • Content Retrieval Pathway (Long Queries, $\ell(q)\geq 150$ characters): Utilizes mean-pooled article embeddings reduced via PCA to 128 dimensions.
  • Headline Retrieval Pathway (Short Queries, $\ell(q)<150$ characters): Utilizes CLS-pooled headline embeddings reduced via PCA to 64 dimensions.
Table XI summarizes the complete configuration of this optimized dual-pathway architecture. The system dynamically routes queries to the appropriate semantic space based on their character length, with each space fine-tuned for maximum relevance through empirically validated pooling strategies and dimensionality reduction.
TABLE XI: Finalized Configuration of the Dual-Pathway Retrieval Pipeline
@@ V-E1 Final Performance Assessment
To validate the effectiveness of the finalized system, we conducted a comprehensive performance assessment using a large-scale test set of 100 queries for each pathway. All true positives (relevant recommendations) were determined through manual evaluation by human assessors, following the rigorous methodology established in Section IV-A. The results, measured by Precision@15, demonstrate the system’s capability to deliver highly accurate recommendations.
@@
For the headline retrieval pathway—designed for concise, keyword-focused queries typical of search bar inputs—the system achieved a mean Precision@15 of 94.35% for 100 queries. The distribution of precision scores across the 100 test queries was highly concentrated, with most queries achieving perfect precision (1.0). This performance represents a significant achievement, as short queries pose a fundamental challenge for semantic matching systems due to their limited contextual information. The 94.35% precision validates our core hypothesis: aligning short queries with semantically compatible headline embeddings effectively bridges the granularity gap that plagues traditional content-based approaches.
@@
For the content retrieval pathway—serving detailed, context-rich queries or full article texts—the system achieved an exceptional mean Precision@15 of 98.53%. While this result is impressive, it aligns with expectations based on our experimental observations. Longer queries inherently contain richer semantic information, enabling more accurate embedding generation and subsequent similarity matching. The high precision achieved here validates the effectiveness of our optimized pipeline, combining mean pooling with PCA reduction to 128 dimensions, for preserving and matching comprehensive semantic contexts. The near-perfect performance across a diverse set of 100 test queries demonstrates the robustness and reliability of this pathway for applications such as ”read-next” article recommendations.
@@ V-E2 Discussion of Results
The differential performance between pathways (94.35% for short queries vs. 98.53% for long queries) reflects the inherent difficulty of each recommendation task. The substantial improvement for short queries, from the initial 60% precision (in case of 13 test set queries) with content embeddings to 94.35% (in case of 100 test set queries) with headline embeddings, represents the primary contribution of our dual-embedding architecture. This transformation from unsatisfactory to production-grade performance addresses a critical limitation in semantic recommendation systems.
@@
The exceptional precision for long queries, while expected given richer input context, nevertheless validates the optimization of each component in the pipeline: appropriate pooling, effective dimensionality reduction, and efficient vector search. The results across both pathways confirm that our systematic experimentation and component-wise optimization have yielded a robust, high-performance recommendation system capable of handling the full spectrum of user query types, from brief keyword searches to detailed contextual inputs.
VI Comparative Analysis with Baseline
Our experimental results demonstrate that the ULTRA framework achieves a Precision@15 of 0.9435 for short, keyword-based 100 queries routed through the headline embedding pathway. This performance is analyzed within the context of the existing landscape for Urdu content recommendation systems, which remains nascent.
VI-A Comparison with an Identified Baseline
A direct, albeit unpublished, baseline for Urdu news recommendation [16] employs a classical TF-IDF approach with cosine similarity. Their methodology involved preprocessing a corpus of approximately 1,160 articles categorized into Business, Science, Entertainment, and Sports. Their reported precision was 0.7143. A comparative summary is presented in Table XII.
@@ VI-A Comparison with an Identified Baseline
TABLE XII: Comparison of the ULTRA Framework with a Baseline Urdu Recommender System
Aspect Baseline System [16] ULTRA Framework (This Work)
Core Methodology TF-IDF with Cosine Similarity Dual-Embedding Architecture (Headline CLS-pooling for $\ell(q)<\theta$; Content Mean-pooling for $\ell(q)\geq\theta$) + PCA
Reported Precision 0.7143 0.9435 (Precision@15 for short queries)
Evaluation Corpus Size $\sim$1,160 articles $\sim$118,000 articles
Relevance Judgment Not explicitly specified Human-evaluated True/False Positives for 100 test queries
Primary Contribution Applies classical IR and BERT embeddings to Urdu news Solves the semantic granularity mismatch; identifies optimal pooling and PCA-reduced dimensions (64D/128D)
The performance of our framework represents a substantial 32% relative improvement in precision over the identified baseline. This significant gain can be attributed directly to our core innovation: the dual-embedding architecture that resolves the semantic granularity mismatch by routing concise queries to a dedicated headline embedding space, optimized through empirical analysis of pooling strategies and PCA-based dimensionality reduction.
Furthermore, our evaluation was conducted on a corpus two orders of magnitude larger ($\sim$118k vs. $\sim$1.16k articles) and employed human-judged relevance assessments, providing a more robust and statistically reliable validation of the reported performance metrics.
VI-B Positioning within Urdu NLP Research
Our work addresses a distinct and relatively underexplored problem within the broader and active field of Urdu Natural Language Processing (NLP). Significant recent research has focused on and achieved high performance (>>92% accuracy/F1-score) in areas such as text classification for news categories, abstractive text summarization using transformer models like mT5, and large-scale fake news detection.
Prior analysis have noted that Urdu information retrieval systems have often relied on foundational techniques like TF-IDF and face challenges due to the language’s morphological complexity and resource scarcity. By successfully designing, optimizing, and validating a transformer-based semantic architecture specifically for content recommendation, this work provides a novel and high-performance pathway for advanced content discovery in Urdu, complementing the advancements in other NLP tasks.
VII Conclusions and Future Research Directions
@@
This research presented and validated the ULTRA framework, a novel dual-embedding architecture designed to address the critical semantic granularity mismatch in Urdu news recommendation. Our core contribution is a system that dynamically routes queries based on character length: concise, keyword-driven queries ($\ell(q)<150$) are matched against CLS-pooled headline embeddings, while detailed, context-rich queries ($\ell(q)\geq 150$) are processed using mean-pooled content embeddings.
The experimental results conclusively demonstrate the efficacy of this approach. The most significant achievement is elevating the precision for short queries, a historically challenging task, from an unsatisfactory $\sim$60% using a single embedding space to a robust 94.35% (Precision@15) by aligning them with semantically compatible headline embeddings and evaluating for a larger test set queries. For long queries, the optimized pipeline achieved an exceptional 98.53% precision, validating the chosen combination of mean pooling and PCA-based dimensionality reduction to 128 dimensions. The systematic evaluation of pooling strategies and dimensionality reduction techniques established PCA as the optimal method for preserving semantic neighborhoods in both embedding spaces, with 64D and 128D identified as the respective optimal dimensions for headline and content pathways.
This work fills a notable gap in Urdu Natural Language Processing (NLP), where prior research has predominantly focused on foundational tasks like text classification and summarization. By developing and rigorously evaluating a sophisticated semantic recommendation system, we provide a validated, high-performance framework for Urdu content discovery.
@@
The success of this framework opens several promising avenues for future work, both to address its limitations and to extend its capabilities. The most direct future direction is the deployment of this system as a production-grade application. A logical next step is to integrate the ULTRA framework into a dedicated Urdu news portal or aggregator. This application could feature a clean search interface for short queries and provide “Read Next” or “Related Articles” panels at the end of articles, powered by the long-query pathway. Moreover, the recommendation engine could be packaged as a cloud-based API service. This would allow existing Urdu digital publishers, blogging platforms, or digital libraries to integrate sophisticated semantic recommendation capabilities into their websites without developing the core NLP infrastructure in-house. The system can also be evolved into a hybrid recommender. The current semantic component could be combined with a collaborative filtering module that analyzes anonymized user interaction data. The final recommendation score could be a weighted combination of semantic similarity and personalized user relevance. Another interesting direction could be figuring out the same dual-embedding logic be applied to recommend related English articles for an Urdu query, or vice-versa, by using a multilingual embedding model. Furthermore, with the rise of multimedia news, could the system be extended to recommend relevant images or short video clips based on the semantic content of an article.

### v2

@@ V-D2 Optimal Dimension for Short Queries: Headline Embeddings
The results, presented in Figure  24, reveal a distinct optimum for the headline embedding space. The 64-dimensional PCA-reduced embeddings achieved the highest overlap score of 63.1%, surpassing both 128D (62.4%) and 256D (61.4%).
Figure 24: Vertical bar chart showing average overlap percentages for different PCA reduced dimensions (64D, 128D, 256D) using headline embeddings
This outcome indicates that for the semantically dense and focused space of headlines, a lower-dimensional representation (64D) is not only sufficient but optimal. Effective semantic capture in a smaller vector space is possible due to the concise nature of headlines where core meaning is compressed in fewer tokens. Extending the dimension to 128D or 256D does not offer any benefit and in fact, has a small negative impact on overlap, probably because it adds redundant or noisy dimensions that do not add any significant variance to this particular data distribution.
Therefore 64D was chosen as the best reduced dimension in headline-based retrieval. This decision is consistent with matching the complexity of representations to the complexity of data: simpler, more narrow semantics of headlines can be well represented in a lower-dimensional space, where they can be stored and accessed efficiently, without losing the fidelity needed to do accurate short-query recommendations.
@@ V-E Final System Configuration and Performance Evaluation
According to the detailed experimental analysis of the results given in the previous sections, the final retrieval pipeline of the ULTRA framework is as follows:
  • Content Retrieval Pathway (Long Queries, $\ell(q)\geq 150$ characters): Embeddings of articles in a mean-pooled representation, downsampled to 128 dimensions with PCA.
  • Headline Retrieval Pathway (Short Queries, $\ell(q)<150$ characters): This uses CLS-pooled headline embeddings reduced to 64 dimensions with PCA.
The entire optimized dual-pathway architecture is summarized in Table IX. The system actively directs queries to the relevant semantic space depending on query length, and each space is optimally relevant using empirically-validated pooling strategies and dimensionality reduction.
TABLE IX: Finalized Configuration of the Dual-Pathway Retrieval Pipeline
@@ V-E1 Final Performance Assessment
The ultimate system was tested through a detailed performance evaluation with an extensive test set of 100 queries per pathway to validate the system effectiveness. The true positives (relevant recommendations) were identified using human evaluation, and it was done manually under the strict methodology, as developed in Section IV-A. The quantitative measures of Precision@15 show that the system is capable of providing very precise recommendations.
@@
Using the headline retrieval pathway which targets brief, keyword-based queries, common in search bars, the system had a mean Precision@15 of 94.35% on 100 queries. The scores of precision on the 100 test queries were concentrated, with the majority of the test queries having a perfect precision (1.0). This performance is a major success because short queries are a major challenge to semantic matching systems since they are not comprehensive in context. The 94.35% precision confirms our fundamental hypothesis: matching short queries to semantically compatible headline embeddings is an effective way of bridging the granularity gap that afflicts traditional content-based methods.
@@
In the case of the content retrieval pathway with detailed, context-rich queries or entire article texts, the system registered an outstanding mean Precision@15 of 98.53%. Although this finding is noteworthy, it is consistent with our experimental findings. The longer queries have more semantic content. Due to this fact, they result in better embedding generation and further similarity matching. The high value of precision confirms that our optimized pipeline, consisting of mean pooling and 128-dimensional PCA reduced embeddings is effective at preserving and matching rich semantic contexts. This performance on a wide range of 100 test queries indicates the strength and stability of this pathway to be implemented in applications like ”read-next” article recommendations.
@@ V-E2 Discussion of Results
The difference in the performance between the pathways (94.35% for short queries and 98.53% for long queries) indicates the difficulty of each recommendation task. The significant enhancement in the short query case using content embeddings, to 94.35% (in case of 100 test set queries) using headline embeddings, is the main contribution of our dual-embedding architecture. This change between unacceptable to production-quality performance addresses a serious weakness in semantic recommendation systems.
@@
The unmatched precision in the long queries, although in the presence of richer input context, however confirms the optimization of all the pipeline components: proper pooling, effective dimensionality reduction, and efficient search of vectors. The outcomes in both directions are an affirmation of the fact that our systematic experimentation and optimization of components has created a robust, high-performance system of recommendation that can support the full range of user query types, both short-term keyword searches to more complex contextual input.
V-F Comparative Analysis with Baselines
It is difficult to create a direct baseline of Urdu content recommendation since the sphere is still young. However, some of the key studies have been the pioneers, and we compare our ULTRA framework with them.
The most applicable published baseline is SEEUNRS (Semantically Enriched Entity-Based Urdu News Recommendation System) [24]. SEEUNRS proposed a two-modular model that included a Semantically Enriched Representation Extractor (SERE) and a Named Entity Extractor (NEE) that used both contextual representations and named entities (persons, places, organizations) to propose relevant articles. On 23,250 news articles curated, the system was tested using the feedback provided by 100 users. Their findings showed that F1 measure was improved by 6.9% compared to traditional methods including kNN, SVD, and TF-IDF. Although SEEUNRS does not provide a direct precision score, this F1 increase provides a very good semantic-enrichment baseline.
Earlier foundational work consists of the Urdu Wikification framework (2022) [17], which created an entity-linking pipeline to recognize the named entities and connect them to Wikidata. This paper referenced 16,738 news stories and built a sub-knowledge graph of 8,439 entities. The system combined the TransE-based knowledge graph embeddings with an RNN to attain a recommendation accuracy of 60.8%. This was the initial evidence of knowledge-enhanced recommendation on Urdu, yet the authors state that the performance is still dependent on the presence of entities in the text.
The other applicable effort is the TF-IDF with BERT embeddings methodology [16], which used the traditional information retrieval methods on a set of about 1,160 Urdu news articles. With cosine similarity, the system achieved a precision of 0.7143 on news recommendation, with the similarity enhanced with BERT embeddings. Nonetheless, this work is based on a small dataset and fails to consider the variability of query length or semantic granularity.
In a wider context, new developments in the area of Urdu document retrieval have created a good performance standard. A 2026 structure (U-RR²) combining TF-IDF, Word2Vec, and mBERT produced P@10 scores of 0.79 - 0.82 and F1 scores of 0.89 - 0.90 with feature ensemble methods on benchmark collections [53]. Although the results in this case refer to document retrieval as opposed to personalized recommendation, the results demonstrate the feasibility of sophisticated Urdu text processing and provide the validation of possible performance levels.
In contrast, our ULTRA framework has a Precision@15 of 0.9435 on short, keyword based queries, which is a significant improvement over previously reported values. This is due to our new dual-embedding architecture with length-threshold routing, which has never been thought in any other Urdu recommendation literature. Moreover, a more formidable validation is delivered by our test on a corpus of around 112,000 articles (orders of magnitude bigger than the 23,250 articles in SEEUNRS) with human-rated relevance scores. Table X summarizes the major differences between ULTRA and other baselines.
@@ V-F Comparative Analysis with Baselines
TABLE X: Comparison of ULTRA with Urdu Recommendation and Retrieval Systems
System Core Methodology Reported Performance Corpus / Dataset
TF-IDF + BERT (2022) [16] TF-IDF + cosine similarity with BERT embeddings Precision = 0.7143 $\sim$1,160 articles (Custom Urdu News)
Urdu Wikification (2022) [17] Entity linking + RNN with TransE knowledge graph embeddings 60.8% accuracy 16,738 articles (SKRS dataset)
SEEUNRS (2024) [24] Entity-based semantic extraction (SERE + NEE) with traditional classifiers 6.9% F1 improvement over TF-IDF/kNN/SVD 23,250 articles (100 users)
U-RR² (2026) [53] Two-stage: (1) TF-IDF + dense embeddings (Word2Vec, FastText, mBERT) + IR models (VSM, BM25, DFR); (2) SVMrank re-ranking with feature ensemble Weighted Word2Vec: MAP 0.78–0.81, P@10 0.79–0.82; Feature Ensemble: F1 0.89–0.90 CURE, ROSHNI, UIR_21 benchmarks
ULTRA (This Work) Dual-Embedding Architecture (Headline CLS-pooling for short queries; Content Mean-pooling for long queries) + PCA Precision@15 = 0.9435 for short queries $\sim$118,000 articles
VI Conclusions and Future Research Directions
@@
The current study introduced and validated the ULTRA framework, a new dual-embedding architecture model that seeks to resolve the problem of severe semantic granularity discrepancy in Urdu content ranking. We make our core contribution with a system that dynamically routes queries by length: short, keyword-based queries ($\ell(q)<150$) are addressed by CLS-pooled headline embeddings, and long, context-rich queries ($\ell(q)\geq 150$) are handled by mean-pooled content embeddings.
The effectiveness of such approach is conclusively demonstrated by the experimental results. The greatest accomplishment has been to bring the precision of short queries, which is traditionally a difficult task, from an unacceptable and unsatisfactory level to a sturdy 94.35% (Precision@15) with an evaluation on a larger test set. In the case of long queries, the optimized pipeline scored a remarkable 98.53% precision, which complimented the selected combination of mean pooling and PCA-based dimensionality reduction to 128 dimensions. The comparative analysis of pooling techniques and dimensionality reduction strategies defined PCA as the best way to retain semantic neighborhoods in both embedding spaces with 64D and 128D as the optimal dimensions of headlines and content paths, respectively.
The work addresses a significant gap in the study of Urdu Natural Language Processing (NLP), in which previous studies have mainly dealt with text classification and summarization as foundational tasks. We offer a validated, high performance content discovery architecture of Urdu by creating and rigorously testing an advanced semantic recommendation system.
The effectiveness of this framework provides multiple avenues of future work, not only to solve its shortcomings but also to enlarge its possibilities. The nearest future move is the implementation of this system as a commercial application. The next logical step is to incorporate the ULTRA framework into a specific Urdu news portal or aggregator. This app may have a simple search interface to support brief searches and offer the long-query pathway-driven ”Read Next” or ”Related Articles” panels at the bottom of articles.
@@
Moreover, the recommendation engine could be packaged as a cloud-based API service. This would allow existing Urdu digital publishers, blogging platforms, or digital libraries to integrate sophisticated semantic recommendation capabilities into their websites without developing the core NLP infrastructure in-house. The system can also be evolved into a hybrid recommender. The current semantic component could be combined with a collaborative filtering module that analyzes anonymized user interaction data. The final recommendation score could be a weighted combination of semantic similarity and personalized user relevance. Another interesting direction could be figuring out the same dual-embedding logic be applied to recommend related English articles for an Urdu query, or vice-versa, by using a multilingual embedding model. Furthermore, with the rise of multimedia news, could the system be extended to recommend relevant images or short video clips based on the semantic content of an article.